OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -4, an = an-1 + 7

OpenStudy (anonymous):

-4, 3, 10, 17, 24, 31 -4, 7, 14, 21, 28, 35 3, 10, 17, 24, 31, 38 0, 7, 14, 21, 28, 35

OpenStudy (anonymous):

@ParthKohli @Hero @.Sam.

OpenStudy (anonymous):

an = an-1 + 7 plug in 2 for n what do you get?

OpenStudy (anonymous):

\[a_{n}=a_{n-1}+7\]

OpenStudy (anonymous):

so whats the next step? @surdawi

OpenStudy (anonymous):

\[n_{2}=n_{2-1}+7\] \[n_{2}=n_{1}+7\] \[n_{2}=(-4)+7\] \[n_{2}=3\] now you try the next 1

OpenStudy (anonymous):

10

OpenStudy (anonymous):

show me how bro